two forces F1 and F2 act on a body of mass 12-kg. F1=7iN and F2=12jN. What are the magnitude and direction relative to the +x axis?
Answer #1

Fnet = F1 + F2

= 7i + 12 j

|Fnet| = sqrt(7^2 + 12^2)

= 13.892 N

a = Fnet/m

= 13.892/12

= 1.158 m/s^2

direction : theta = tan^-1(Fy/Fx)

= tan^-1(12/7)

= 59.74 degrees above +x axis